OpenStudy (anonymous):

Solve for x: 1 over 4 (2x − 14) = 4.

OpenStudy (michele_laino):

is your equation this? \[\frac{ 1 }{ 4(2x-14) }=4\]

OpenStudy (anonymous):

x:1/4 (2x − 14) = 4 is what it had on the paper, i don't understand what it wants, honestly

OpenStudy (michele_laino):

Ok! I try

OpenStudy (anonymous):

these are answer choices if that helps any 15 15/2 9 11

OpenStudy (michele_laino):

I found!

OpenStudy (michele_laino):

here are all steps: your equation is: \[\frac{ 1 }{ 4 }(2x-14)=4\] so multiply both sides by 4, we get: \[2x-14=16\] then adding 14 to both sides, we can wite: \[2x=30\]so dividing both sides by 2 we have the solution: \[x=15\]

OpenStudy (anonymous):

Wow, thanks man, I think I understand now.

OpenStudy (michele_laino):

thank you!